OpenStudy (anonymous):

what is 40% of 3/4 in decimal form and fraction form?

OpenStudy (anonymous):

steps please

OpenStudy (anonymous):

30

OpenStudy (anonymous):

.4 x .75

OpenStudy (anonymous):

3/4 X 40/100 = 120/400 = 3/10

OpenStudy (anonymous):

as a decimal = .3

OpenStudy (anonymous):

.5 x .75=.3